OpenStudy (anonymous):

A 0.50 kg ball moves in a circle 40 cm in radius at a speed of 4.0 m/s. (a) Find it's centripetal acceleration. (b) What is the value of the centripetal force on the ball?

OpenStudy (anonymous):

Centripetal acceleration is given as follows.\[\boxed{\displaystyle a_c = \frac{v^2}{r}}\]By Newton's second law \(F=ma\), we can find the centripetal force.\[F_c=ma_c \Rightarrow \boxed{\displaystyle F_c=\frac{mv^2}{r}}\]